OpenStudy (anonymous):

I'm looking at trying to develop an absolute value function to model the path of a ball in a pool game. The ball begins at (o,1) and is aimed at the topside of the table at y=10. It must then bounce back and must strike the second ball at (8,3). The function must be in the form y=a|x+b|+c. I am unsure of how to develop a function with only two given points. Thank-you!

pooja195 (pooja195):

@triciaal

pooja195 (pooja195):



pooja195 (pooja195):

we have three points (0,1) , (x,10) ( 8,3) lets see if we can plug them into y = a| x + b | + c 1 = a | 0 + b | + c 10 = a | x + b | + c 3 = a | 8 + b | + c

pooja195 (pooja195):

And you went offline...

OpenStudy (anonymous):

hi sorry i am coming back I had an assignment with partners and had to meet them

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

Hi sorry what is the next step. Do I rearrange and then sub each into each other to find A B C- I am not sure how to work with 3 variables. Can you help with that ? Thanks so much.

hartnn (hartnn):

those are 3 equations with 4 unknowns, those are insufficient, if we can find x', we will have all 3 points,  so you know how to get the equation of line given 2 points? get the equation of line with points (0,1) and (x',10) get the equation of line with points (3,8) and (x',10) solving them simultaneously can get us the value of x'. thinking of a simpler approach...